# SOC-DGL: Social Interaction Behavior Inspired Dual Graph Learning Framework for Drug-Target Interaction Identification

### v1

@@ Abstract
The identification of drug-target interactions (DTI) is crucial for drug discovery and repositioning, as it reveals potential uses of existing drugs, aiding in the acceleration of the drug development process and reducing associated costs. Despite the similarity information in DTI is important, most models are limited to mining direct similarity information within homogeneous graphs, overlooking the potential yet rich similarity information in heterogeneous graphs. Inspired by real-world social interaction behaviors, we propose SOC-DGL, which comprises two specialized modules: the Affinity-Driven Graph Learning (ADGL) module and the Equilibrium-Driven Graph Learning (EDGL) module. The ADGL module adopts a comprehensive social interaction strategy, leveraging an affinity-enhanced global drug-target graph to learn both global DTI and the individual similarity information of drugs and targets. In contrast, the EDGL module employs a higher-order social interaction strategy, amplifying the influence of even-hop neighbors through an even-polynomial graph filter grounded in balance theory, enabling the indirect mining of higher-order homogeneous information. This dual approach enables SOC-DGL to effectively and comprehensively capture similarity information across diverse interaction scales within the affinity matrices and drug-target association matrices, significantly enhancing the model’s generalization capability and predictive accuracy in DTI tasks. To address the issue of imbalance in drug-target interaction datasets, this paper proposes an adjustable imbalance loss function that mitigates the impact of sample imbalance by adjusting the weight of negative samples and a parameter. Extensive experiments on four benchmark datasets demonstrate significant accuracy improvements achieved by SOC-DGL, particularly in scenarios involving data imbalance and unseen drugs or targets.
@@
Drug-target interaction, affinity-driven graph learning, equilibrium-driven graph learning, data imbalance.
@@ 1 Introduction
Drug repurposing, the strategy of applying approved medications to novel therapeutic indications, has gained prominence due to the stagnation in traditional drug discovery [1]. Drug-target interactions (DTI) are central to pharmacology, yet conventional drug discovery remains time-consuming, costly, and high-risk [2]. In contrast, repurposing seeks to identify new uses for existing drugs, significantly reducing the time and cost of drug development [3]. Consequently, computational models for drug-target interactions have emerged, utilizing previous experimental data to predict novel DTI strengths. With the expanding availability of biomedical data, artificial intelligence-based methods for drug repurposing have proliferated  [4] , primarily falling into two key categories: machine learning-based methods and increasingly popular network-based methods.
Predicting DTI typically involves mapping drug-target pairs to feature vectors and applying machine learning for classification [5, 6, 7]. However, machine learning-based methods are limited by the need for positive and negative samples and are susceptible to sample imbalance, which can introduce noise and reduce prediction accuracy. Network-based methods typically build homogeneous or heterogeneous biological networks, infer interactions from drug-drug similarity, target-target similarity, or DTI, and predict interactions via network reasoning [8, 9, 10, 11]. Although network-based reasoning methods are effective to some extent, they primarily focus on direct interactions and fail to explore similarity information embedded in heterogeneous information networks.
To address this challenge, this study proposes a novel social interaction-inspired dual-enhanced graph reasoning mechanism, termed SOC-DGL, which integrates ADGL and EDGL for accurate and robust DTI prediction based on similarity information from diverse sources. The main contributions of this study are:
  • Inspired by real-world social interactions, this study proposes two complementary modules, ADGL and EDGL. Based on ADGL, which is built upon an extensive social interaction framework to achieve global representation learning on affinity-enhanced drug-target networks, EDGL further incorporates balance theory and employs an even-polynomial graph filter to capture high-order even-hop neighbor relationships in the heterogeneous drug-target graph. These two modules work in a complementary manner, enabling a comprehensive exploration of homogeneous information for both drugs and targets across multiple interaction scales.
  • To mitigate the impact of imbalanced data on the performance of SOC-DGL, this study designs a novel an adjustable imbalance loss function that mitigates the impact of sample imbalance by weighting negative samples and introducing an adjustable parameter $\varpi$. Extensive experimental results on imbalanced benchmark datasets consistently demonstrate the effectiveness.
@@ 2.1 DTI Prediction Based on Similarity
The integration of similarity information plays a pivotal role in advancing DTI prediction models. Traditionally, similarity mining has focused on isomorphic networks, where pairwise similarity between drugs and targets is utilized [12, 13, 14]. However, recent work has shifted towards heterogeneous graphs, which integrate multiple types of nodes and relationships [15, 16, 17]. Despite these advancements, these approaches primarily rely on similarity information derived from traditional similarity matrices, overlooking the rich and potentially valuable indirect similarity information between drugs and targets embedded within heterogeneous graphs. Such higher-order similarity information is critical for predicting potential DTI associations and necessitates the design of specialized modules for its learning.
@@ 2.2 High-order Similarity Information Learning
Effectively learning high-order similarity information aids in uncovering complex higher-order isomorphic relationships in graph data. Mi et al.[18] proposed HSL, which learns high-order similarity by considering both first-order and high-order similarities in a low-dimensional embedding space. Wang et al.[19] introduced HSDA for speech emotion recognition, capturing high-order similarity through a similarity graph to reveal cross-domain structural relationships. Peng et al. [20] proposed HSC, where high-order similarity is learned from relationships between dimensional spaces and jointly optimized with ordinary similarity for clustering. Inspired by these methods and balance theory[21], we propose the EDGL module, built upon the ADGL module, to fully exploit high-order similarity information for DTI prediction.
@@ 3 Datasets and Methodology
3.1 Datasets
To comprehensively assess the performance and applicability of SOC-DGL, this study selects four representative benchmark datasets: KIBA [22], Davis [23], BindingDB [24], and DrugBank [25], whose detailed description appears in Appendix Section A.
3.2 Model framework overview
The framework of SOC-DGL is composed of six main steps, as shown in Fig. 1, including: (I) Affinity-Enhanced Global Drug-Target Network Construction, (II) Affinity-Driven Graph Learning, (III) Equilibrium-Driven Graph Learning, (IV) Feature Fusion, (V) DTI Prediction Decoder, and (VI) Imbalanced Processing.
@@ 3.2.1 Affinity-Enhanced Global Drug-Target Network Construction
In DTI prediction, diverse features of drugs and targets are derived from various measurement methods, offering complementary perspectives on molecular characteristics. Integrating these heterogeneous features is crucial for accurate DTI inference. Assuming a shared latent space for drug and target features, SOC-DGL employs pairwise multi-view learning to derive a unified subspace representation from multiple views, specifically the drug affinity matrix ($A_{DD}$) and target affinity matrix ($A_{TT}$). SOC-DGL effectively integrates multi-view information by constructing a protocol matrix that combines correlation matrices under low-rank and sparsity constraints. Details are provided in Appendix Section B.
@@ 3.2.2 Affinity-Driven Graph Learning
To extract both direct and global DTI and similarity information of drugs and targets, this study employs an extensive social interaction strategy and apply global graph convolutional learning to the affinity-enhanced global drug-target network matrix.
To effectively filter the useful information within the heterogeneous graph, for drug affinity matrix $A_{DD}$ and the target affinity matrix $A_{TT}$, elements in these two affinity matrices that are greater than or equal to the threshold 0.8 are set to 1, and the rest are set to 0, resulting in the drug affinity adjacency matrix $\tilde{A}_{DD}$ and the target affinity adjacency matrix $\tilde{A}_{TT}$. By combining the drug-target affinity matrix, this study constructs the adjacency matrix $G$ of:
@@
Subsequently, this study performs graph convolution operations:
@@
where $D^{-1/2}GD^{-1/2}$ represents the normalization process, with $D=\mathrm{diag}\left(\sum_{j}G_{ij}\right)$ being the degree matrix of the adjacency matrix $G$. $H^{(l)}$ is the feature embedding matrix at layer $l$, $W^{(l)}$ is the learnable weight matrix at layer $l$, and $\sigma$ is the activation function. After two rounds of graph convolution, the graph convolution feature embedding matrix $H^{\prime}$ is obtained:
@@ 3.2.3 Equilibrium-Driven Graph Learning
In the real situation, drugs with similar structures may exhibit analogous mechanisms of action, while the similarity between targets reflects the functional conservation of targets [26]. As the classic adage goes: “The friend of my friend is also my friend,” this concept is equally applicable to DTI. Using balance theory [21], this study deduces that even though drugs and targets differ in their attributes, they can establish connections through shared interaction relationships. If a drug interacts with multiple targets, and these targets are connected to other drugs, then these drugs and targets may form a state of equilibrium through these indirect interaction relationships. To this end, this study proposes the EDGL module, which employs a higher-order social interaction strategy to indirectly learn higher-order drug and target similarity information linked by ’friend nodes’ through an even-polynomial graph filter.
This study processes $H^{\prime}$, which has been encoded through ADGL. For ease of expression, let $X_{0}=H^{\prime}$, and then perform two linear transformations on $X_{0}$, applying an activation function. $X$ is obtained as follows:
@@
To learn similarity information in drug-target association matrix, this study sets $(\tilde{A}_{DD}^{\prime})_{ij}=0$, $(\tilde{A}_{TT}^{\prime})_{ij}=0$, thereby transforming it into the adjacency matrix $A^{\prime}$:
$A^{\prime}=\begin{pmatrix}0&A_{DT}^{\prime}\\ A_{TD}^{\prime}&0\end{pmatrix}\in\mathbb{R}^{(n_{d}+n_{t})\times(n_{d}+n_{t})}$ (7)
where $A_{DP}^{\prime}\in\mathbb{R}^{n_{d}\times n_{t}}$, $\tilde{A}_{DD}^{\prime}\in\mathbb{R}^{n_{d}\times n_{d}}$. This study then obtains the Laplacian matrix $L=D^{\prime}-A^{\prime}$, where $D^{\prime}$ is the degree matrix of $L$. After standardizing $L$, the formula is as follows:
@@
where $I$ is the identity matrix. Subsequently, this study defines the propagation matrix as $P=I-\tilde{L}=D^{-1/2}A^{\prime}D^{-1/2}$, and the $k$-th order polynomial graph filter is defined as $g(\tilde{L})=\sum_{k=0}^{K}\alpha_{k}\tilde{L}^{k}$, $k=0,1,\ldots,K$, where $\alpha_{k}$ is the learnable weight matrix, and $\alpha_{k}=\alpha\times(1-\alpha)^{k}$, with $\alpha$ being a hyperparameter that can be set. Thereafter, this study rewrites the graph filter formula as $g(\tilde{L})=\sum_{k=0}^{K}\alpha_{k}(I-\tilde{L})^{k}=\sum_{k=0}^{K}\alpha_{k}P^{k}$, discarding the odd-degree terms in $g(\tilde{L})$ and retaining only even-degree terms:
$g_{\text{even}}(\tilde{L})=\sum_{k=0}^{\left\lfloor\frac{K}{2}\right\rfloor}\alpha_{2k}(I-\tilde{L})^{2k}=\sum_{k=0}^{\left\lfloor\frac{K}{2}\right\rfloor}\alpha_{k}P^{2k}$ (9)
Thus, the model of EDGL is constructed:
$Y=f(\hat{X},P)=\sum_{k=0}^{\left\lfloor\frac{K}{2}\right\rfloor}\alpha_{k}P^{2k}\hat{X}$ (10)
where $\hat{X}$ is the input feature matrix that has already undergone two linear transformations, and $Y$ represents the final output features. This study denotes the result obtained after passing $H^{\prime}$ through EDGL as $H^{\prime\prime}$, and its expression is as follows:
$H^{\prime\prime}=Y=\begin{bmatrix}H_{D}^{\prime\prime}\\ H_{T}^{\prime\prime}\end{bmatrix}\in\mathbb{R}^{(n_{d}+n_{t})\times m}$ (11)
@@ 3.2.4 Feature Fusion
This study employs residual connection techniques to integrate information from diverse sources, ensuring that the two modules complement each other and collectively form a more powerful graph reasoning mechanism. ”This approach balances the influence of higher-order interaction patterns on the generalized patterns, achieving equilibrium between the two ”social interaction paradigms.” The corresponding formula is as follows:
$\hat{H}=\omega H^{\prime}+(1-\omega)H^{\prime\prime}$ (12)
@@ 3.2.5 DTI prediction decoder
DTI prediction decoder in SOC-DGL cleverly employs the principle of matrix factorization. By meticulously decoding these features, it reveals the affinity between drugs and targets, assessing the potential interaction affinity between different drugs and targets. This study denotes $W^{L}$ as the learnable weight matrix for the mapping from the drug space to the target space within the decoder. Inspired by matrix tri-factorization, the formula for DTI prediction decoder is as follows:
$H^{*}=\sigma(\hat{H}_{D}W^{(L)}\hat{H}_{P}^{T})$ (13)
where $H^{*}$ represents the final predicted interaction matrix, and its element $H_{ij}^{*}$ denotes the predicted association score between drug $i$ and target $j$.

### v2

@@ Abstract
The identification of drug-target interactions (DTI) is critical for drug discovery and repositioning, as it reveals potential therapeutic uses of existing drugs, accelerating development and reducing costs. However, most existing models focus only on direct similarity in homogeneous graphs, failing to exploit the rich similarity in heterogeneous graphs. To address this gap, inspired by real-world social interaction behaviors, we propose SOC-DGL, which comprises two specialized modules: the Affinity-Driven Graph Learning (ADGL) module, learning global similarity through an affinity-enhanced drug-target graph, and the Equilibrium-Driven Graph Learning (EDGL) module, capturing higher-order similarity by amplifying the influence of even-hop neighbors using an even-polynomial graph filter based on balance theory. This dual approach enables SOC-DGL to effectively capture similarity information across multiple interaction scales within affinity and association matrices. To address the issue of imbalance in DTI datasets, we propose an adjustable imbalance loss function that adjusts the weight of negative samples by the parameter. Extensive experiments on four benchmark datasets demonstrate that SOC-DGL consistently outperforms existing state-of-the-art methods across both balanced and imbalanced scenarios. Moreover, SOC-DGL successfully predicts the top 9 drugs known to bind ABL1, and further analyzed the 10th drug, which has not been experimentally confirmed to interact with ABL1, providing supporting evidence for its potential binding.
@@
drug-target interaction, dual graph learning, social interaction, data imbalance.
@@ 1 Introduction
Drug repurposing, the strategy of applying approved medications to novel therapeutic indications, has gained prominence due to the stagnation in traditional drug discovery [1]. Drug-target interactions (DTI) are central to pharmacology, yet conventional drug discovery remains time-consuming, costly, and high-risk compared to repurposing [2]. Consequently, computational methods have emerged to predict novel DTI strengths. With the exploding of biomedical data, artificial intelligence-based methods for drug repurposing have proliferated [3], including two key categories: machine learning-based methods and network-based methods.
Predicting DTI typically involves mapping drug-target pairs to feature vectors and applying machine learning for classification [4]. However, these machine learning-based methods are limited by the need for positive and negative samples and are susceptible to sample imbalance, which can introduce noise and reduce prediction accuracy. Neural networks have been widely applied in various fields [5, 6, 7, 8, 9, 10, 11]. Network-based methods typically build homogeneous or heterogeneous biological networks, infer interactions from drug-drug similarity, target-target similarity, or DTI [12]. Although network-based methods are effective to some extent, they primarily focus on direct interactions and fail to explore similarity information embedded in heterogeneous information networks.
Most existing DTI prediction methods primarily focus on mining direct similarity information within homogeneous graphs, neglecting the potentially rich high-order similarity information embedded in heterogeneous drug-target networks [13]. This oversight limits the ability to fully capture complex and indirect relationships essential for accurate DTI identification, especially under data imbalance scenarios. Therefore, there is a clear research gap in designing effective models that can comprehensively exploit both direct and high-order similarity information from heterogeneous graphs.
To address this challenge, this study proposes a novel social interaction-inspired dual-enhanced graph reasoning mechanism, termed SOC-DGL, whose main contributions are:
  • Inspired by real-world social interactions, this study proposes two complementary modules, ADGL and EDGL, enabling a comprehensive exploration of homogeneous information for both drugs and targets across multiple interaction scales. Based on ADGL, which is built upon an extensive social interaction framework to achieve global representation learning on affinity-enhanced drug-target networks, EDGL further incorporates balance theory and employs an even-polynomial graph filter to capture high-order even-hop neighbor relationships in the heterogeneous drug-target graph.
  • To mitigate the impact of imbalanced data on SOC-DGL, this study designs an adjustable imbalance loss function that weights negative samples and introduces an adjustable parameter $\varpi$. Extensive experimental results on imbalanced benchmark datasets consistently demonstrate the effectiveness.
@@
The objective of this study is to explore rich similarity information from heterogeneous drug-target graphs by a dual graph learning framework inspired by social interaction strategies, thereby enhancing the accuracy of DTI prediction. The rest of this article is organized as follows. Section II reviews related work to SOC-DGL. Section III presents the datasets and details the SOC-DGL framework. Section IV describes the experimental settings and discusses results. Finally, Section V concludes the paper and outlines future research directions.
@@ 2 Related Work
SOC-DGL is closely tied to previous studies for DTI prediction, particularly those related to DTI prediction based on similarity and high-order similarity information learning.
@@ 2.1 DTI Prediction Based on Similarity
The integration of similarity information plays a pivotal role in advancing DTI prediction models. Traditionally, similarity mining has focused on isomorphic networks, where pairwise similarity between drugs and targets is utilized. DTiGEMS+ [14], combines graph embedding, mining, and similarity techniques for improved drug–target interaction prediction. iNGNN-DTI [15], an interpretable graph neural network, integrates pre-trained models and enhances accuracy while offering insights into DTI. However, recent work has shifted towards heterogeneous graphs, which integrate multiple types of relationships. Shao et al. introduced DTI-HETA [16], an end-to-end model that integrates heterogeneous graphs and attention mechanisms for improved DTI prediction. Notably, GIAE-DTI [17] combines cross-modal information from drugs and targets with self-supervised learning, demonstrating promising performance. However, these methods primarily rely on information derived from traditional similarity matrices, overlooking the rich and potentially valuable indirect similarity information between drugs and targets embedded within heterogeneous graphs. Such higher-order similarity information is critical for predicting potential DTI associations.
@@ 2.2 High-order Similarity Information Learning
Effectively learning high-order similarity information aids in uncovering complex higher-order isomorphic relationships in graph data. In HSC [18], high-order similarity is learned from relationships between dimensional spaces and jointly optimized with ordinary similarity for clustering. HSL [19] learns high-order similarity by considering both first-order and high-order similarities in a low-dimensional embedding space. DisCo [20], a graph-based disentangled contrastive learning framework captures high-order user similarity relationships through multi-step random walks. scHNTL [21] identifies subtle cell similarities by constructing high-order adjacency matrices and optimizing cell embeddings using triplet loss. HSDA [22] is introduced for speech emotion recognition, capturing high-order similarity through a similarity graph to reveal cross-domain structural relationships. Inspired by these methods and balance theory [23], this study propose the EDGL module, built upon the ADGL module, to fully exploit high-order similarity information for DTI prediction.
@@ 3 Datasets and Methodology
In reality, drugs with similar structures may exhibit analogous mechanisms of action, while the similarity between targets reflects the functional conservation of targets [24]. Based on the observation that in human social networking, people engage in both direct friendships and higher-order connections through mutual friends, we draw an analogy in DTI prediction. We model the ”direct friendships” by learning the known associations between drugs and targets. For the ”higher-order friendships,” we employ an even-polynomial graph filter based on balance theory. According to the triadic stability principle in balance theory, in a triadic relationship, if two relationships are positive, the third relationship tends to be positive as well [23]. For instance, aspirin and salicylic acid share similar chemical structures and both exert their effects by inhibiting cyclooxygenase (COX) [25]. We hypothesize that, due to the underlying commonality between social interactions and drug-target relationships, strategies inspired by social friendship patterns can be effectively applied to DTI prediction [26].
In this paper, we propose a dual-graph learning framework that simultaneously captures both direct similarities and higher-order indirect relationships within the heterogeneous drug-target graph, enhancing the accuracy of DTI prediction.
3.1 Datasets and Preprocessing
To comprehensively assess the performance and applicability of SOC-DGL, this study selects four representative benchmark datasets: KIBA [27], Davis [28], BindingDB [29], and DrugBank [30], whose detailed description appears in Appendix Section A. For each benchmark dataset, we extract various features of drugs and targets based on the SMILES sequences of drugs and the UniProt IDs of targets. For drug sequences, we use the RDKit software package to extract Molecular Access System (MACCS) keys fingerprints [31], topological (TOP) fingerprints [32], and Morgan fingerprints [33], which represent three distinct aspects of the molecular features. For target sequences, to provide a more systematic description of target sequences and their physicochemical properties, we employ the iLearn software package to calculate amino acid composition (AAC), Moran autocorrelation (MORAN) [25], Composition/Transition/Distribution (CTD) [34], and pseudo-amino acid composition (PAAC) [35], which encompass four different aspects of target features.
3.2 Methodology
The framework of SOC-DGL is as shown in Fig. 1.
@@ 3.2.1 Affinity-Enhanced Global Drug-Target Network Construction
In DTI prediction, diverse features of drugs and targets are derived from various measurement methods, offering complementary perspectives on molecular characteristics. Integrating these features is crucial for accurate DTI inference. Assuming a shared latent space for drug and target features, SOC-DGL employs pairwise multi-view learning to derive a unified subspace representation from multiple views, specifically the drug affinity matrix ($A_{DD}$) and target affinity matrix ($A_{TT}$). SOC-DGL effectively integrates multi-view information by constructing a protocol matrix that combines correlation matrices under low-rank and sparsity constraints. Details are provided in Appendix Section B (1).
@@ 3.2.2 Affinity-Driven Graph Learning
To effectively filter the useful information within the heterogeneous graph, for $A_{DD}$ and $A_{TT}$, elements in these two affinity matrices that are greater than or equal to the threshold 0.8 are set to 1, and the rest are set to 0, resulting in the drug affinity adjacency matrix $\tilde{A}_{DD}$ and the target affinity adjacency matrix $\tilde{A}_{TT}$. Therefore, the adjacency matrix $G$ is constructed as:
@@
Then, this study performs graph convolution operations:
@@
where $D^{-1/2}GD^{-1/2}$ is the normalization process, with $D=\mathrm{diag}\left(\sum_{j}G_{ij}\right)$ being the degree matrix of $G$. $H^{(l)}$ and $W^{(l)}$ are the feature embedding matrix and learnable weight matrix at layer $l$. $\sigma$ is the activation function. After two rounds of graph convolution, the feature embedding matrix $H^{\prime}$ is obtained:
@@ 3.2.3 Equilibrium-Driven Graph Learning
For ease of expression, let $X_{0}=H^{\prime}$, and perform two linear transformations, applying an activation function. $X$ is obtained as follows:
@@
To learn similarity information in drug-target association matrix, this study sets $\tilde{A}_{DD}^{\prime}=0$, $\tilde{A}_{TT}^{\prime}=0$, thereby transforming it into the adjacency matrix $A^{\prime}$. Then, this study obtains the Laplacian matrix $L=D^{\prime}-A^{\prime}$, where $D^{\prime}$ is the degree matrix of $L$. After standardizing $L$, $\tilde{L}$ is obtained:
@@
where $I$ is the identity matrix. Subsequently, this study defines the propagation matrix as $P=I-\tilde{L}=D’^{-1/2}A^{\prime}D^{\prime-1/2}$, and the $k$-th order polynomial graph filter is defined as $g(\tilde{L})=\sum_{k=0}^{K}\alpha_{k}\tilde{L}^{k}$, $k=0,1,\ldots,K$, where $\alpha_{k}$ is the learnable weight matrix, and $\alpha_{k}=\alpha\times(1-\alpha)^{k}$, with $\alpha$ being a hyperparameter that can be set. Thereafter, this study rewrites the graph filter formula as $g(\tilde{L})=\sum_{k=0}^{K}\alpha_{k}(I-\tilde{L})^{k}=\sum_{k=0}^{K}\alpha_{k}P^{k}$, discarding the odd-degree terms in $g(\tilde{L})$ and retaining only even-degree terms:
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}H^{\prime\prime}=\sum_{k=0}^{\left\lfloor\frac{K}{2}\right\rfloor}\alpha_{k}P^{2k}{X}}$ (8)
where $X$ is the input feature matrix that has already undergone two linear transformations.
@@ 3.2.4 Feature Fusion
This study employs residual connection techniques to integrate information from diverse sources, ensuring that the two modules complement each other and collectively form a more powerful graph reasoning mechanism. This approach balances the influence of higher-order interaction patterns on the generalized patterns, achieving equilibrium between the two social interaction paradigms. The corresponding formula is provided in Appendix Section B (2).
@@ 3.2.5 DTI Prediction Decoder
DTI prediction decoder in SOC-DGL cleverly employs the principle of matrix factorization. By meticulously decoding these features, it reveals the affinity between drugs and targets, assessing the potential interaction affinity between different drugs and targets. This study denotes $W^{L}$ as the learnable weight matrix for the mapping from the drug space to the target space within the decoder. Inspired by matrix tri-factorization, the formula is as follows:
$H^{*}=\sigma(\hat{H}_{D}W^{(L)}\hat{H}_{T}^{T})$ (9)
where $\hat{H}_{D}$ and $\hat{H}_{T}$ are obtained through feature fusion, $H^{*}$ represents the final predicted interaction matrix.